# Conceptualization, Data Curation Conceptualization Visualization, Writing - Review

### v1

@@ 4.6 Transformers and Self-attention Networks
A summary of various algorithms that were successfully used for condition monitoring of TEP, along with their specifications is presented in Table 3.
| Model | Reference | Citation | Specification | Faults Information | Average Accuracy |
LSTM |
[114] |
Yuan 2019 |
@@
Discrete wavelet transform and CNN for feature extraction and softmax classifier |
1, 2, 4, 8, 10, 12, 13, 14, 17, and 18 |
98.73% |
BiRNN |
[120] |
Zhang 2019 |
@@
Bidirectional recurrent neural networks using gated recurrent unit (GRU) |
All |
92.70% |
LSTM + CNN |
[76] |
Shao 2019 |
@@
First LSTM and then CNN are used for feature extraction. The detection of faults 3, 9, and 15 are improved. |
All |
92.06% |
RNN |
[36] |
Kang 2020 |
Using a 4 layer RNN, however, could not detect fault 15 |
All |
92.50% |
LSTM |
[36] |
Kang 2020 |
Better performance compared to RNN speciallly for fault 15 |
All |
95% |
LSTM |
[118] |
Zhang 2020 |
Three layers LSTM are stacked |
All |
99.49% |
BiRNN |
[13] |
Chadha 2020 |
Bidirectional recurrent neural networks |
All |
88.14% |
GRU |
[56] |
Mirzaei 2022 |
The detetcion of fault 15 is enhanced compared to LSTM |
All |
95.90% |
CNN |
[92] |
Wu 2018 |
CNN with dropout |
20 faults |
88.20% |
CNN |
[116] |
Zhang 2019 |
FFT for feature extraction |
All faults except 3,9,15,16 |
96.90% |
CNN |
[21] |
Gao 2020 |
@@
Handling multi time scale data using fixed multiple sampling method |
All |
98.53% |
CNN |
[28] |
He 2021 |
@@
Multiblock temporal convolutional network (MBTCN), dividing process variables into subblocks |
All faults except 3,9,15 |
96.20% |
CNN |
[111] |
Yu 2021 |
1D CNN and Wavelet transform for feature extraction |
All |
94.08% |
CNN |
[86] |
Wang 2021 |
Normalized CNN to improve the stability and speed of the training |
All |
92.75% |
CNN |
[26] |
Guo 2022 |
@@
Multiblock adaptive convolution kernel neural network (MBCKN) for large scale process using attention mechanism |
All |
94.15% |
Autoencoder |
[35] |
Jin 2018 |
Using wavelet activation function in the autoencoder |
All |
95.86% |
Autoencoder |
[108] |
Yu 2020 |
@@
Convolutional LSTM memory autoencoder (CLSTM-AE) and residual block for increasing the accuracy and make the training process more robust |
All |
91% |
Autoencoder |
[110] |
Yu 2020 |
@@
Manifold regularized stacked autoencoders (MRSAE) considering neighbourhood information needs a few labelled data |
All |
91.42% |
Autoencoder |
[121] |
Zhang 2022 |
@@
Vector autoregressive model and 1-dimensional convolutional autoencoder for capturing the process dynamics |
All faults except 3, 9 and 15 |
96.80% |
Autoencoder |
[60] |
Pan 2021 |
@@
Neuron-grouped stacked autoencoder (CG-SAE) as a hierarchical feature extractor for different groups of faults |
All faults except3, 9 and 15 |
96.02% |
Autoencoder |
[10] |
Bi 2021 |
@@
Orthogonal self-attentive variational autoencoder (OSAVA) to increase the interpretability |
All faults except 15, 16 |
94.70% |
Autoencoder |
[38] |
Li 2021 |
@@
Slow feature analysis-aided autoencoder (SFA-AE) to deal with slow variations in the data |
All faults except15, 17 |
94.87% |
Autoencoder |
[1] |
Agarwal 2022 |
Combining LSTM and Supervised Autoencoder |
All |
93.23% |
Autoencoder |
[15] |
Cheng 2019 |
Variational recurrent autoencoder combines RNN and VAE |
All faults except 3,9,15 |
96.30% |
Autoencoder |
[106] |
Yu 2022 |
@@
Supervised convolutional autoencoder (SCAE) for better feature extraction |
All faults except 3,9,15,16 |
96.19% |
Autoencoder |
[85] |
Wang 2023 |
Convolutional adversarial autoencoder (1DAAE) |
All |
86% |
Autoencoder |
[101] |
Yang 2023 |
@@
Nonlocal and global preserving stacked autoencoder (NLGPSAE) using a regularized objective function |
All |
94.80% |
DBN |
[88] |
Wang 2020 |
@@
Extended deep belief network (EDBN)enhances the data quality during feature compression |
All faults except3,9,15 |
94.31% |
DBN |
[102] |
Yao 2021 |
Adaptive lifting wavelet-DBN (ALW-DBN) to reduce the noise level |
All |
97.33% |
DBN |
[100] |
Yang 2021 |
@@
Pairwise graph regularized deep belief network (PG-DBN) to monitor and control the feature manifold structure of the data |
All |
93.87% |
Transformer |
[117] |
Zhang 2022 |
Self attention transformer |
All |
80.10% |
Transformer |
[90] |
Wei 2022 |
@@
Target transformer using both self-attention and target-attention mechanism |
20 faults |
90.39% |
Transformer |
[129] |
Zhou 2023 |
Vision Transformer and CNN Increasing the accuracy, especially for faults 3,9 and 15 |
20 faults |
94.08% |
Attention mechanism |
[119] |
Zhang 2019 |
@@
Combines self-attention (SA) mechanism to the gated recurrent unit (GRU) |
20 faults |
95.04% |
Attention mechanism |
[96] |
Xiong 2023 |
@@
Attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) with Dropout |
21 faults |
88.80% |
@@
Recurrent Networks |
Process sequential or temporal data, modelling time dependencies |
Prone to gradient vanishing/exploding |
Attention Models |
Handling sequential data and weighting important samples in the input sequence. No fixed window size is needed. Can capture slow changes in long time series and extract features of long-term dependencies |
Needs rich historical data |
CNN |
Powerful feature extraction capability. Has denoising capabilities [47]. |
May need multiple layers and large labelled datasets to perform well |
Autoencoder |
Feature extraction and dimension reduction ability. Unsupervised training. Robust against noise |
Pretraining needed |
GAN |
Generate data to enhance the data imbalance problem. Can be used for semi-supervised classification [72] |
Stability problem in the training |
DBN |
Can be trained unsupervised or supervised. Energy-based probabilistic models |
Computationally expensive training |
@@ 5.3 Imbalanced Data
Zhang et al. [123]proposed a zero-shot learning (ZSL) approach to handle unseen faults. Zero-shot learning involves the utilization of expert-provided high-level descriptions instead of relying on trained objects to identify target items. These descriptions encompass semantic attributes such as colour, shape, and behaviours, which can be learned beforehand without requiring samples from unfamiliar categories. In essence, ZSL is an approach to train models to recognize patterns in new types of data using information gleaned from familiar classes along with relevant descriptions. At first feature extraction from raw data is done using a one-dimensional convolutional neural network (1D CNN), and then labelling is done by semantic descriptions by expert knowledge. Five Different subsets of TEP are used to validate the proposed method, considering different faults as unseen data. The detection accuracies for different scenarios vary from 59.72% to 96.67%, which outperforms other ZSL methods like SAE in most of the scenarios.
@@
Guo et al. [25], used a modified stacked autoencoder to learn low-level features containing the information of complete data based on incomplete input data. This method creates data representations that can be used for classification. The steps of the data driven fault detection process are shown in 11. When the data is not rich, two general approaches can be handle the problem. Augmenting or generating more data, or using the models that can perform with few samples. Different scenarios and possible solutions are shown in the mentioned figure. Table 5 shows the summary of methods that can handle imbalanced, incomplete or unlabeled data.
Method |
Reference |
Citation |
Specification |
Fault Information |
Average Accuracy |
Unlabelled |
[57] |
Morales 2019 |
@@
Combination of autoencoder and LSTM model and a control chart for anomaly detection of unlabelled data |
6.25% of the data is labelled |
97.46% |
Unlabelled |
[67] |
Peng 2020 |
@@
Bidirectional gated recurrent unit (BGRU), weighting training example to overcome the class imbalance and cost-sensitive active learning to explore the unlabelled data |
Data imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 |
97.50% |
Unlabelled and Imbalanced |
[65] |
Peng 2020 |
@@
Enhanced auxiliary classifier generative adversarial network (EACGAN) to deal with imbalanced data and more stable training |
Normal to fault ratio of 10:1, all faults except 3, 9 and 15 |
77.01% |
Imbalanced |
[128] |
Zheng 2022 |
@@
Three-step high-fidelity positive-unlabeled (THPU) is used to deal with having a few labels in the data set, by generating and augmenting data |
One labelled normal sample is available the rest is unlabelled |
92.32% |
Unlabelled and Imbalanced |
[63] |
Peng 2022 |
@@
Contrastive learning for information extraction of unlabeled and imbalanced data |
Only 20 samples of each fault is available |
86.59% |
Imbalanced |
[4] |
Amini 2022 |
@@
Source-aware autoencoder (SAAE) for handling imbalanced data combining BiLSTM and ResNet |
Detecting unseen faults |
93.80% |
Unlabelled |
[51] |
Ma 2022 |
@@
Consistency regularization autoencoder (CRAE) for handling unlabelled data |
All faults except 3, 9, and 15 (20% of data is labelled) |
92.95% |
Unlabelled and Imbalanced |
[66] |
Peng 2022 |
@@
Using genetic algorithm and sparse autoencoder for imbalanced and unlabeled fault diagnosis |
Unlabelled data and the normal to fault ratio of 10:1 |
92.02% |
Imbalanced |
[72] |
Qin 2022 |
@@
High-efficiency GAN using Wasserstein GAN and Auxiliary Classifier, to generate and classify data for imbalanced dataset |
Normal to fault ratio of 5:1 |
82.00% |
Unlabelled |
[122] |
Zhang 2022 |
@@
Combined LSTM for temporal feature extraction and ladder autoencoder (LAE) to handle unlabeled data |
All faults except 3,13, 15 and 16, 11% of the dataset is labelled |
95.90% |
Unknown Faults |
[55] |
Mirzaei 2023 |
@@
Few-shot learning and GRU with attention layer for unknown fault identification |
Faults 0, 2, 12, and 14 as known faults, Faults 1, 4, 5, 7, 8, 10, 11, and 13 as unknown |
85.70% |
Imbalanced |
[97] |
Xu 2023 |
@@
Few-shot learning based on a class-rebalance strategy to detect new faults. Feature extraction is done using 1D-CNN with an attention module |
3 to 50 samples of faults for each class |
From 81.8% to 94.7% |
Imbalanced |
[52] |
Ma 2023 |
@@
Double branch rebalanced network (DBRN) to deal with imbalance classes using resampling and reweighting |
Imbalance ratio is 10:1 for faults 1,2,6 |
98.75% |

### v2

@@ 4.6 Transformers and Self-attention Networks
A summary of various algorithms that were successfully used for condition monitoring of TEP, along with their specifications is presented in Table LABEL:tab:models.
Model | Reference | Specification | Faults Information | Average Accuracy |
@@
LSTM | [118] | Discrete wavelet transform and CNN for feature extraction and softmax classifier | 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18 | 98.73% |
@@
BiRNN | [124] | Bidirectional recurrent neural networks using gated recurrent unit (GRU) | All | 92.70% |
@@
LSTM + CNN | [79] | First LSTM and then CNN are used for feature extraction. The detection of faults 3, 9, and 15 are improved. | All | 92.06% |
RNN | [38] | Using a 4 layer RNN, however, could not detect fault 15 | All | 92.50% |
LSTM | [38] | Better performance compared to RNN speciallly for fault 15 | All | 95% |
LSTM | [122] | Three layers LSTM are stacked | All | 99.49% |
BiRNN | [13] | Bidirectional recurrent neural networks | All | 88.14% |
GRU | [58] | The detetcion of fault 15 is enhanced compared to LSTM | All | 95.90% |
CNN | [96] | CNN with dropout | 20 faults | 88.20% |
CNN | [120] | FFT for feature extraction | All faults except 3,9,15,16 | 96.90% |
@@
CNN | [23] | Handling multi time scale data using fixed multiple sampling method | All | 98.53% |
@@
CNN | [30] | Multiblock temporal convolutional network (MBTCN), dividing process variables into subblocks | All faults except 3,9,15 | 96.20% |
CNN | [115] | 1D CNN and Wavelet transform for feature extraction | All | 94.08% |
CNN | [90] | Normalized CNN to improve the stability and speed of the training | All | 92.75% |
@@
CNN | [28] | Multiblock adaptive convolution kernel neural network (MBCKN) for large scale process using attention mechanism | All | 94.15% |
CNN + LSTM | [75] | CNN is used to extract multi-level fault features followed by atcked LSTM solving the gradient disappearance problem. | All | 95.19% |
CNN + LSTM | [131] | Extracts multiscale features for various operating conditions | All faults except 3,9,15,21 | 98% |
CNN + LSTM | [15] | Improved feature extraction and prevent model degradation | All faults | 98% |
Autoencoder | [37] | Using wavelet activation function in the autoencoder | All | 95.86% |
@@
Autoencoder | [112] | Convolutional LSTM memory autoencoder (CLSTM-AE) and residual block for increasing the accuracy and make the training process more robust | All | 91% |
@@
Autoencoder | [114] | Manifold regularized stacked autoencoders (MRSAE) considering neighbourhood information needs a few labelled data | All | 91.42% |
@@
Autoencoder | [125] | Vector autoregressive model and 1-dimensional convolutional autoencoder for capturing the process dynamics | All faults except 3, 9 and 15 | 96.80% |
@@
Autoencoder | [62] | Neuron-grouped stacked autoencoder (CG-SAE) as a hierarchical feature extractor for different groups of faults | All faults except 3, 9 and 15 | 96.02% |
@@
Autoencoder | [10] | Orthogonal self-attentive variational autoencoder (OSAVA) to increase the interpretability | All faults except 15, 16 | 94.70% |
@@
Autoencoder | [40] | Slow feature analysis-aided autoencoder (SFA-AE) to deal with slow variations in the data | All faults except 15, 17 | 94.87% |
Autoencoder | [1] | Combining LSTM and Supervised Autoencoder | All | 93.23% |
Autoencoder | [17] | Variational recurrent autoencoder combines RNN and VAE | All faults except 3,9,15 | 96.30% |
@@
Autoencoder | [110] | Supervised convolutional autoencoder (SCAE) for better feature extraction | All faults except 3,9,15,16 | 96.19% |
Autoencoder | [89] | Convolutional adversarial autoencoder (1DAAE) | All | 86% |
@@
Autoencoder | [105] | Nonlocal and global preserving stacked autoencoder (NLGPSAE) using a regularized objective function | All | 94.80% |
@@
DBN | [92] | Extended deep belief network (EDBN) enhances the data quality during feature compression | All faults except 3,9,15 | 94.31% |
DBN | [106] | Adaptive lifting wavelet-DBN (ALW-DBN) to reduce the noise level | All | 97.33% |
@@
DBN | [104] | Pairwise graph regularized deep belief network (PG-DBN) to monitor and control the feature manifold structure of the data | All | 93.87% |
Transformer | [121] | Self attention transformer | All | 80.10% |
@@
Transformer | [94] | Target transformer using both self-attention and target-attention mechanism | 20 faults | 90.39% |
Transformer | [134] | Vision Transformer and CNN increasing the accuracy, especially for faults 3,9 and 15 | 20 faults | 94.08% |
@@
Attention mechanism | [123] | Combines self-attention (SA) mechanism to the gated recurrent unit (GRU) | 20 faults | 95.04% |
@@
Attention mechanism | [100] | Attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) with Dropout | 21 faults | 88.80% |
@@
Recurrent Networks | Process sequential or temporal data, modelling time dependencies | Prone to gradient vanishing/exploding |
Attention Models | Handling sequential data and weighting important samples in the input sequence. No fixed window size is needed. Can capture slow changes in long time series and extract features of long-term dependencies | Needs rich historical data |
CNN | Powerful feature extraction capability. Has denoising capabilities [49]. | May need multiple layers and large labelled datasets to perform well |
Autoencoder | Feature extraction and dimension reduction ability. Unsupervised training. Robust against noise | Pretraining needed |
GAN | Generate data to enhance the data imbalance problem. Can be used for semi-supervised classification [74] | Stability problem in the training |
DBN | Can be trained unsupervised or supervised. Energy-based probabilistic models | Computationally expensive training |
@@ 5.1 Unlabeled Data
Current zero-shot learning (ZSL) methods mostly focus on global level fault features; however, local level fault features that characterize the topological relationships between neighboring data points, are very important for generalization, to solve this problem a global–local attentionaware ZSL (GLA-ZSL) method is proposed in [85] consisting of a global prior refinement module using 2D-CNN for global features and local feature enhancement module based on the channel attention and 1D-CN, in addition a meta-learner is used for generalization. Human-defined fault attributes are used in this study.
@@ 5.3 Imbalanced Data
Zhang et al. [127]proposed a ZSL approach to handle unseen faults. Zero-shot learning involves the utilization of expert-provided high-level descriptions instead of relying on trained objects to identify target items. These descriptions encompass semantic attributes such as colour, shape, and behaviours, which can be learned beforehand without requiring samples from unfamiliar categories. In essence, ZSL is an approach to train models to recognize patterns in new types of data using information gleaned from familiar classes along with relevant descriptions. At first feature extraction from raw data is done using a one-dimensional convolutional neural network (1D CNN), and then labelling is done by semantic descriptions by expert knowledge. Five Different subsets of TEP are used to validate the proposed method, considering different faults as unseen data. The detection accuracies for different scenarios vary from 59.72% to 96.67%, which outperforms other ZSL methods like SAE in most of the scenarios.
@@
Guo et al. [27], used a modified stacked autoencoder to learn low-level features containing the information of complete data based on incomplete input data. This method creates data representations that can be used for classification. The steps of the data driven fault detection process are shown in 11. When the data is not rich, two general approaches can be handle the problem. Augmenting or generating more data, or using the models that can perform with few samples. Different scenarios and possible solutions are shown in the mentioned figure. Table LABEL:tab:Imb shows the summary of methods that can handle imbalanced, incomplete or unlabeled data.
Method | Reference | Specification | Fault Information | Average Accuracy |
@@
Unlabelled | [59] | Combination of autoencoder and LSTM model and a control chart for anomaly detection of unlabelled data | 6.25% of the data is labelled | 97.46% |
@@
Unlabelled | [69] | Bidirectional gated recurrent unit (BGRU), weighting training example to overcome the class imbalance and cost-sensitive active learning to explore the unlabelled data | Data imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 | 97.50% |
@@
Unlabelled and Imbalanced | [67] | Enhanced auxiliary classifier generative adversarial network (EACGAN) to deal with imbalanced data and more stable training | Normal to fault ratio of 10:1, all faults except 3, 9 and 15 | 77.01% |
@@
Imbalanced | [133] | Three-step high-fidelity positive-unlabeled (THPU) is used to deal with having a few labels in the data set, by generating and augmenting data | One labelled normal sample is available the rest is unlabelled | 92.32% |
@@
Unlabelled and Imbalanced | [65] | Contrastive learning for information extraction of unlabeled and imbalanced data | Only 20 samples of each fault is available | 86.59% |
@@
Imbalanced | [4] | Source-aware autoencoder (SAAE) for handling imbalanced data combining BiLSTM and ResNet | Detecting unseen faults | 93.80% |
@@
Unlabelled | [53] | Consistency regularization autoencoder (CRAE) for handling unlabelled data | All faults except 3, 9, and 15 (20% of data is labelled) | 92.95% |
@@
Unlabelled and Imbalanced | [68] | Using genetic algorithm and sparse autoencoder for imbalanced and unlabeled fault diagnosis | Unlabelled data and the normal to fault ratio of 10:1 | 92.02% |
@@
Imbalanced | [74] | High-efficiency GAN using Wasserstein GAN and Auxiliary Classifier, to generate and classify data for imbalanced dataset | Normal to fault ratio of 5:1 | 82.00% |
@@
Unlabelled | [126] | Combined LSTM for temporal feature extraction and ladder autoencoder (LAE) to handle unlabeled data | All faults except 3,13, 15 and 16, 11% of the dataset is labelled | 95.90% |
@@
Unknown Faults | [57] | Few-shot learning and GRU with attention layer for unknown fault identification | Faults 0, 2, 12, and 14 as known faults, Faults 1, 4, 5, 7, 8, 10, 11, and 13 as unknown | 85.70% |
@@
Imbalanced | [101] | Few-shot learning based on a class-rebalance strategy to detect new faults. Feature extraction is done using 1D-CNN with an attention module | 3 to 50 samples of faults for each class | From 81.8% to 94.7% |
@@
Imbalanced | [54] | Double branch rebalanced network (DBRN) to deal with imbalance classes using resampling and reweighting | Imbalance ratio is 10:1 for faults 1,2,6 | 98.75% |